# Empirical investigation of multi-source cross-validation in clinical machine learning

### v1

@@ 3.2 ECG classifier
All data were resampled to 250 Hz and each ECG was adjusted to a length of 4096 samples, approximately 16 seconds. Signals over 16 seconds were randomly cropped and zero padding was randomly applied to shorter signals. The signal amplitudes were normalized to the range [0,1]. Additionally, gender values were encoded using one-hot encoding, and age values were scaled to the range [0,1].
@@ 3.3.1 Classifier performance metrics
The metrics are computed for a filtered subset of the log probabilities provided by the Sigmoid function and the corresponding actual labels. This means, before computation, we filtered out all the labels and their corresponding predictions that do not occur within the actual labels. For example, as shown in Table 3, the PTB and PTB-XL dataset do not contain any ECGs indicating right bundle branch block, and therefore, we filtered them out to facilitate the computation of the metrics.
@@ 3.3.2 Cross-validation methods
In this study, we consider two variants of cross-validation (CV): stratified K-fold cross-validation (K-fold CV), and leave-source-out CV (LSO CV). Figure 2 illustrates the difference between the two CV methods. In CV, the dataset is repeatedly partitioned into a training set used for fitting a model, and validation set used to evaluate the prediction performance of the model. The performance estimates from different CV repetitions are averaged to obtain the final performance estimate. Compared to using a single data split, CV has the advantage of allowing the use of all the data to validate the model. CV methods have also been used derive confidence intervals for prediction performance, though this has been shown to be challenging due to dependencies between the repetitions [35].
Figure 2: Illustration of the difference between the K-fold CV and the LSO CV: In the K-fold CV, a training set is first pooled and then divided into K disjoint folds, whereas in the LSO CV, the data sources correspond to distinct folds.
@@
Standard K-fold CV consists of randomly partitioning the dataset to K equally sized non-overlapping folds (typically, K=10 or K=5) [36]. On each round of CV, K-1 folds are use as the training set, and the remaining fold as the validation set. In stratified K-fold CV the folds are sampled in a way that preserves the class distribution in the original dataset. Stratification can be especially important when classes with very low membership are present in the data, to ensure that sufficient representation from each class in present in all the folds. K-fold CV can be expected to perform well if the training set can be assumed to be an i.i.d. sample from the same population on which the prediction model will later be applied on. This basic assumption is likely to be violated in applications where the goal is to make predictions on data from new sources (e.g. a hospital) not represented in the data used to develop the model. This can lead to unreliable performance estimation.
@@
LSO CV [21, 10] is a method designed to be applied on data arising from different sources which have different generating processes. The goal is to estimate the out-of-source prediction performance. The method works by partitioning the data so that each fold corresponds to data points coming from a single source. On each round of CV, one source is left out as validation set, and the rest are combined to form the training set. In the context of our study, each source corresponds to a single data source described e.g. in Table 1, and LSO CV estimates the expected prediction performance of the model for patients from a new source not part of the data used to develop and validate the model.
@@ 3.3.3 Performance metrics for comparing cross-validation methods
To measure how well K-fold CV and LSO CV estimate prediction performance on a new source, the following metrics are employed. We perform repeated experiments on $n$ distinct data sources, each corresponding to patients from one source. Let $i\in\{1,2,...,n\}$ be the index for an out-of-source test dataset, $\hat{y}_{i}$ the CV AUC and $y_{i}$ the test AUC for a model trained on data from other sources. Further, let $e_{i}=\hat{y}_{i}-y_{i}$ denote the signed error for the CV estimate. Then, $\mu=\frac{1}{n}\sum_{i=1}^{n}e_{i}$ denotes the mean error (ME) over the experiments, $\sigma=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(e_{i}-\mu)^{2}}$ the standard deviation (SD), and $\sqrt{\mu^{2}+\sigma^{2}}$ the root mean square error (RMSE). The ME captures any systematic bias, with positive value corresponding to overoptimistic and negative to pessimistic CV results. The standard deviation measures the random component of the error, and together these form the overall RMSE.
3.4 Experimental setting
@@
The experiments were run using Python software package 3.10.4 with Torch version 1.13.1. We used the Puhti supercomputer provided by the CSC – IT Center for Science Ltd. to run and evaluate the Se-ResNet models. CSC offers high-quality information technology services to academic institutions, research institutes and businesses in Finland. The code is freely available on GitHub, with the link provided in the Code Availability section.
@@ 3.4.1 Experimental setup for single source data
In the single source experiments, we repeated the same procedure for each pair of sources in the data. First, one source was chosen as the training and other as the test set. Then, stratified 5-fold CV estimate for AUC was calculated on the training set. Then we trained a model on the full training set, and evaluated the AUC on the test set. The ME, SD, and RMSE was calculated for each training source separately, by comparing the 5-fold CV AUC to the test AUCs of the model on the other sources. One model was trained on each fold, so in total, 25 models were trained for calculating the CV estimates, and five models for final testing.
@@ 3.4.2 Experimental setup for multi-source data
In the multi-source experiments, one data source was consistently left out as a test set, and the remaining four sources were used together as a training set. For the training set, we employed two CV methods: the stratified 4-fold CV, where folds consist of pooled data from all the sources, as well as the LSO CV, where data from each source corresponded to its own fold (Figure 2). The reason for choosing the value of K as 4 is because LSO also divides the training set into 4 folds, with each fold corresponding to one data source. One model was trained per one fold, resulting in 20 (= the number of CV folds multiplied by the number of data sources) models per CV method, and therefore, 40 models in total.
Finally, each training set was used to fit the final model and the test set was used to evaluate the performance. Thus, we ended up training five final models, one for each test set. The ME, SD, and RMSE were computed for both the CV methods over five experiments, each corresponding to a different choice of test source.
@@ 4 Results
We carried out two distinct sets of experiments. These experiments were designed to provide insight on the reliability of CV estimates, when the final goal is to apply the trained model to patient populations from new data sources not represented in the training data.
@@ 4.1 Single source experiments
The results for the first experiments, where data from a single source was used for training and CV are summarized in Table 4. The rows in the table correspond to the test sets, the columns to the training set, and the 5-fold CV results are represented on the diagonal.
The CV results are consistently higher than any test result corresponding to the same training set. The model trained on the CPSC and CPSC-Extra dataset had the lowest macro-AUC CV estimate (0.905), but resulted in test AUCs between 0.793–0.857 ($\mu$=0.810). The highest macro-AUC CV estimate was computed for the model trained on the SPH dataset (0.980) with test AUCs ranging between 0.891–0.931 ($\mu$=0.908). Even though the CV estimate for the model trained on the Chapman-Shaoxing and Ningbo dataset was second highest (0.968), the model resulted in highest mean of test AUCs (0.926). For micro-AUC CV, the lowest was for G12EC (0.957) with test AUCs ranging between 0.832 – 0.968 ($\mu$=0.910) and highest for SPH (0.993) with test AUCs ranging between 0.907 – 0.942 ($\mu$=0.926). The lowest mean of the test AUCs was for the model trained on the CPSC and CPSC-Extra dataset (0.808), the CV estimate being relatively high (0.982).
The ME, SD and RMSE for between CV AUCs and test AUCs are summarized in Table 5 per training source. The mean errors are positive for all of the datasets, ranging from 0.015 – 0.096 for macro-AUC and 0.047 – 0.174 for micro-AUC, demonstrating overoptimistic bias for 5-fold CV estimates when compared to test performance on an independent data source.
Table 4: Results for different training and test set combinations, with 5-fold CV results on the diagonal (macro-AUC/micro-AUC).
Train set
Test set
Chapman-Shaoxing
and Ningbo
CPSC and
CPSC-Extra
G12EC
PTB and
PTBXL
SPH
Chapman-Shaoxing
and Ningbo
@@
CPSC and
CPSC-Extra
[…]
PTB and
PTBXL
[…]
Table 5: Assessment of the reliability of the 5-fold CV per training set in the single-source experiment (macro-AUC/micro-AUC)
Chapman-Shaoxing
and Ningbo
CPSC and
CPSC-Extra
G12EC
PTB and
PTBXL
SPH
Mean 0.042 / 0.074 0.096 / 0.174 0.015 / 0.047 0.070 / 0.139 0.072 / 0.068
SD 0.025 / 0.050 0.032 / 0.071 0.029 / 0.057 0.032 / 0.099 0.017 / 0.016
RMSE 0.049 / 0.089 0.101 / 0.188 0.033 / 0.074 0.077 / 0.171 0.074 / 0.070
4.2 Multi-source experiments
The results from the second set of experiments, where data from multiple sources were used for training and validation are summarized in Figure 3. The results are reported per test set. The results of the 4-fold CV were higher than the LSO CV for both macro- and micro-AUC (Figure 3), as expected. The 4-fold CV AUCs are very similar regardless of the test source, and variance over the folds is small. In contrast, there is more variability between LSO CV results for different test sources and also between the folds.
The comparison between the errors of 4-fold CV and LSO CV estimates is presented in Table 6. The mean errors for the LSO CV are close to zero (macro-AUC: -0.0013, micro-AUC: -0.0062), indicating unbiased performance estimates. In contrast, the 4-fold CV results mean errors demonstrate clear optimistic bias (macro: 0.0404, micro: 0.0489). On the other hand, the LSO CV has higher SD for the errors (macro: 0.0326 $>$ 0.0259, micro: 0.0378 $>$ 0.0265). Combining these two components of error together, LSO CV has lower RMSE (macro: 0.0327 $<$ 0.0480, micro: 0.0383 $<$ 0.0556) providing overall more reliable performance estimates in the experiments than 4-fold CV.
[…]
@@ 4.2 Multi-source experiments
4-fold CV LSO CV
Macro-AUC
Micro-AUC
Macro-AUC
Micro-AUC
Mean 0.0404 0.0489 -0.0013 -0.0062
SD 0.0259 0.0265 0.0326 0.0378
RMSE 0.0480 0.0556 0.0327 0.0383

### v2

@@ 3.2 Predictive models
The data preprocessing was identical for both neural networks and followed similar principles as in [8]. All data were resampled to 250 Hz and each ECG was adjusted to a length of 4096 samples, approximately 16 seconds. Signals over 16 seconds were randomly cropped. For shorter signals, zero padding was applied to both sides of the signals to extend them to the required sample size, with the proportion of padding between the sides determined randomly. The signal amplitudes were normalized to the range [0,1]. Additionally, sex values were encoded using one-hot encoding, and age values were scaled to the range [0,1].
For the benchmark models, LR and XGBoost, we extracted 20 features from lead II, replicating the approach used in the transformer implementation [36], which also incorporated these hand-crafted features. They employed a random forest to assess feature importance and selected the top 20 features from over 300 extracted ECG features based on the results. These features include heart rate variability features and morphological features. Additionally, age and sex features were included in the feature set, resulting in a total of 22 features that are used in classification.
3.2.1 Improved residual neural network
The first model [8] is a residual neural network (ResNet) architecture with additional squeeze and excitation (SE) blocks. The ResNet consists of a single convolutional layer, followed by eight residual blocks (RBs). Each of these blocks comprises two convolutional layers and the SE block. Both the first convolutional layer and the RB units employ 64 convolutional filters, with the number of filters doubling after every second RB unit. The feature dimension is reduced by half after the max pooling layer and the third, fifth, and seventh RBs. The ResNet was used with binary cross entropy loss and the Adam optimizer with an initial learning rate of 0.003.
Compared to the original ResNet architecture, the improved ResNet includes four modifications: Firstly, the age and sex information are passed to the network through an additional fully connected layer before being added to the final fully connected layer. Secondly, to enhance the learning of meaningful features, a kernel size of 15 is applied to the first convolutional kernel, while the size of the other kernels is 7. Thirdly, a dropout layer with a dropout rate of 0.2 is placed between the two convolutional kernels in each RB to reduce the risk of overfitting. Finally, a SE block is added into each RB to model the spatial relationships between the ECG channels.
3.2.2 Wide and deep transformer neural network
The second model [36] is a wide and deep transformer neural network. The wide component incorporates the 22 hand-crafted features, while the deep component includes three modules: 1) An embedding network that applies a series of convolution operations to capture the latent space representation of the ECG signal; 2) a transformer module; and 3) a multi-label classification head with fully connected layers to generate 64 deep features. These deep features are concatenated with the wide features and fed into a final fully connected layer that produces outputs corresponding to the diagnostic labels.
Similarly to the ResNet, binary cross entropy loss was used. However, instead of the Adam optimizer, the Noam optimization was employed as in the original implementation. This approach incorporates the Adam optimizer within a scheduling routine that linearly increases the learning rate during a 4000-step warm up phase and then decreases it according to the inverse square root of the step number.
3.2.3 Benchmark models
For the LR classifier, the multilabel classification was performed using one-vs-the-rest strategy, where one classifier is fit for each label. The maximum number of iterations was adjusted to 2000 to ensure convergence. Aside from this, the default hyperparameter values were used. The XGBoost classifier was trained using default hyperparameter values.
@@ 3.3.1 Classifier performance metrics
The metrics are computed for a filtered subset of the predictions, along with the corresponding actual labels. This means, before computation, we filtered out labels and their corresponding predictions that not present in the actual labels of the training or test sets. For example, as shown in Table 3, the PTB and PTB-XL source do not contain any ECGs indicating right bundle branch block, and therefore, we filtered them out to facilitate the computation of the metrics.
@@ 3.3.2 Cross-validation methods
In this study, we consider two variants of cross-validation (CV): stratified K-fold cross-validation (K-fold CV), and leave-source-out CV (LSO CV). Figure 2 illustrates the difference between the two CV methods. In CV, the dataset is repeatedly partitioned into a training set used for fitting a model, and validation set used to evaluate the prediction performance of the model. The performance estimates from different CV repetitions are averaged to obtain the final performance estimate. Compared to using a single data split, CV has the advantage of allowing the use of all the data to validate the model. CV methods have also been used derive confidence intervals for prediction performance, though this has been shown to be challenging due to dependencies between the repetitions [38]. Regardless of CV approach used, the final model is trained on all the training data combined together.
Figure 2: Illustration of the difference between the K-fold CV and the LSO CV. The training data is gathered from $n$ distinct sources. In the K-fold CV, a training set is first pooled from all the sources and then randomly divided into K disjoint folds, whereas in the LSO CV, the data sources correspond to distinct folds. Each split corresponds to one round of CV, where a model is trained on the union of train folds and used to make predictions on the validation fold. The final performance estimate is the average of the validation fold performances. The final model is trained on the full training set consisting of all the data from the $n$ sources. A test source refers to out-of-sample data that is not part of the $n$ sources used for training and cross-validation, but on which the model would later be applied on.
@@
Standard K-fold CV consists of randomly partitioning the dataset to K equally sized non-overlapping folds (typically, K=10 or K=5) [39, 40]. On each round of CV, K-1 folds are use as the training set, and the remaining fold as the validation set. In stratified K-fold CV the folds are sampled in a way that preserves the class distribution in the original dataset. Stratification can be especially important when classes with very low membership are present in the data, to ensure that sufficient representation from each class in present in all the folds. K-fold CV can be expected to perform well if the training set can be assumed to be an i.i.d. sample from the same population on which the prediction model will later be applied on. This basic assumption is likely to be violated in applications where the goal is to make predictions on data from new sources (e.g. a hospital) not represented in the data used to develop the model. This can lead to unreliable performance estimation. For the standard i.i.d. setting, K-fold cross-validation is known to have a negative bias, because models trained on K-1 folds tend to perform slightly worse than ones trained with the full training set, especially for small values of K. Empirically, it has been shown that with large enough K (e.g. K=10 or $20$) this issue is mostly resolved [40]. On the other hand, for multi-source data, K-fold can be expected to have positive bias for out-of-source prediction performance estimation, because same the sources are represented in both training and validation folds [10].
@@
LSO CV [21, 10] is a method designed to be applied on data arising from different sources which have different generating processes. The goal is to estimate the out-of-source prediction performance. The method works by partitioning the data so that each fold corresponds to data points coming from a single source. On each round of CV, one source is left out as validation set, and the rest are combined to form the training set. In the context of our study, each source corresponds to a single data source described e.g. in Table 1, and LSO CV estimates the expected prediction performance of the model for patients from a new source not part of the data used to develop and validate the model. While LSO CV does not suffer from the same positive bias as K-fold does, it can be expected to have a negative bias [10]. This is because it estimates the prediction performance for a model trained on n-1 sources, whereas the final model trained on all the n sources is likely to perform better, especially if the number of sources is small. This negative bias cannot be reduced straightforwardly by increasing the number folds, since this would require access to new data sources.
@@ 3.3.3 Performance metrics for comparing cross-validation methods
To measure how well K-fold CV and LSO CV estimate prediction performance on a new source, the following metrics are employed. We perform repeated experiments on $n$ distinct data sources, each corresponding to patients from one source. Let $i\in\{1,2,...,n\}$ be the index for an out-of-source test source, $\hat{y}_{i}$ the CV AUC and $y_{i}$ the test AUC for a model trained on data from other sources. Further, let $e_{i}=\hat{y}_{i}-y_{i}$ denote the signed error for the CV estimate. Then, $\mu=\frac{1}{n}\sum_{i=1}^{n}e_{i}$ denotes the mean error (ME) over the experiments, $\sigma=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(e_{i}-\mu)^{2}}$ the standard deviation (SD), and $\sqrt{\mu^{2}+\sigma^{2}}$ the root mean square error (RMSE). The ME captures any systematic bias, with positive value corresponding to overoptimistic and negative to pessimistic CV results. The standard deviation measures the random component of the error, and together these form the overall RMSE.
3.4 Experimental setup
@@
The experiments were run using Python software package 3.10.4 with Torch version 1.13.1. We used the Puhti supercomputer provided by the CSC – IT Center for Science Ltd. to run and evaluate the neural network models. CSC offers high-quality information technology services to academic institutions, research institutes and businesses in Finland. The LR classifier was implemented using the scikit-learn Python library version 1.3.1 and the XGBoost classifier was implemented using xgboost Python library version 2.1.0. The code is freely available on GitHub, with the link provided in the Code Availability section.
@@ 3.4.1 Experimental setup for single source data
In the single source experiments, we repeated the same procedure for each pair of sources in the data: First, one source was chosen as a training set and another as a test set. To estimate performance on the training source, the stratified 5-fold CV was conducted. Here, the performance estimate was calculated as the average over the five train/validation folds performed within the CV.
Finally, we trained a model on the full training set, and evaluated the AUC on the test set. The ME, SD, and RMSE were calculated for each training source separately, by comparing the 5-fold CV AUC to the test AUCs of the model on the other sources.
@@ 3.4.2 Experimental setup for multi-source data
In the multi-source experiments, one data source was left out as a test set, and the remaining four sources were used together as a training set. The experiment was repeated five times, so each source served in turn as the test set.
For the training set, we employed two CV methods: the stratified 4-fold CV, where folds consist of pooled data from all the sources, as well as the LSO CV, where data from each source corresponded to its own fold (Figure 2). The reason for choosing the value of K as 4 is because LSO also divides the training set into 4 folds, with each fold corresponding to one data source in the training set.
The full training set was used to fit the final model and the test set was used to evaluate the performance. Thus, we ended up training five final models, one for each test set. The ME, SD, and RMSE were computed for both the CV methods over five experiments, each corresponding to a different choice of test source. The final models for both the 4-fold CV and the LSO CV were the same, as the training set always included four out of five sources, with the remaining source used as the test set.
3.4.3 Experimental setup for source prediction experiment
The central assumption underlying the use of LSO CV is that there are systematic differences between data coming from different sources. From Figure 1 and Table 3 one can already observe the clear differences in label distributions. We further employ a heuristic based on training a multi-class classifier to evaluate the degree to which the data sources can be distinguished from each other. In this experiment, we combined the data from all the sources and divided it with 70 % allocated for training and 30 % for testing sets using stratification based on the sources. A ResNet classifier [8] was then trained to predict, from which of the five data sources a patient comes from. If all the data came from same distribution, the expected classification performance would be on random level, or at most equal to relative fraction of ECGs coming from the largest source for a majority voter classifier. Achieving higher classification performance can be taken as indicative of there being clear differences between the sources.
Three different input sets were used to determine which features contribute most significantly to accurate hospital identification. The input sets were 1) ECGs, sex and age, representing the initial feature set used in the subsequent multilabel classification; 2) ECG, sex, age and diagnoses, i.e. the labels in the ECG classification task; and 3) diagnoses only. The diagnoses were incorporated through the final fully-connected layer, consistent with the original architecture of the ResNet, which integrated age and sex features in this manner. Additionally, for the multiclass classification task where the outputs were the data sources $s_{n},n=1,...,5$, the Softmax function was used for activation instead of the originally used Sigmoid function, and cross entropy was used for computing the loss. We visualized the outcomes in confusion matrices to identify whether some sources were more prone to misclassification and which source would commonly be misclassified as another. For classification metric, we computed classification accuracy.
@@ 4 Results
We carried out two distinct sets of experiments in single source and multi-source settings. These experiments were designed to provide insight on the reliability of CV estimates, when the final goal is to apply the trained model to patient populations from new data sources not represented in the training data. Additionally, we analyzed how accurately one can distinguish between the patient cohorts across the five data sources to gain a deeper understanding of the differences and potential biases among the sources.
@@ 4.1 Single source experiments
The results for the experiments, where data from a single source was used for training and CV are summarized in Figure 3 per training set. It’s noteworthy that the models trained on either of the CinC sources result in the highest test macro- and micro-AUC for the SPH source. In some instances, the test AUCs exceed the corresponding CV estimates.
The CV results are generally higher than test result corresponding to the same training set for each model. In terms of the macro-AUC, the CPSC and CPSC-Extra source show greater variation in CV estimates compared to the other data sources, regardless of which model was trained. The ResNet achieves higher CV and test AUCs than the other models irrespective of the test source, whereas the Transformer tends to exhibit greater variation in CV estimates. A similar trend is observed for the micro-AUC, with larger variations in test AUCs: Regardless of the trained model, the CV estimates often exceed 0.9 and surpass the corresponding test AUCs. When models are trained on the CPSC and CPSC-Extra source, the micro-AUCs are also lower compared to those trained on other data sources.
Figure 3: AUC ($\pm$ SD) scores for the 5-fold CV and individual test sets per training set
The ME, SD and RMSE for between CV AUCs and test AUCs are summarized in Table 4 per training source. With a single exception, the MEs are positive for all the models regardless of the training source, demonstrating overoptimistic bias for 5-fold CV estimates when compared to test performance on an independent data source.
Table 4: Assessment of the reliability of the 5-fold CV per training set in the single-source experiment
Macro-AUC Micro-AUC Train source Method Mean SD RMSE Mean SD RMSE Chapman-Shaoxing and Ningbo ResNet 0.0434 0.0260 0.0506 0.0754 0.0518 0.0914 Transformer 0.1338 0.0477 0.1420 0.1703 0.1207 0.2088 LR 0.0618 0.0318 0.0695 0.2383 0.0838 0.2526 XGBoost 0.1040 0.0234 0.1066 0.1730 0.0704 0.1868 CPSC and CPSC-Extra ResNet 0.0586 0.0276 0.0648 0.1544 0.0738 0.1711 Transformer 0.0529 0.0304 0.0610 0.2156 0.1185 0.2460 LR 0.0535 0.0220 0.0579 0.2068 0.0869 0.2244 XGBoost 0.0358 0.0234 0.0428 0.1465 0.0747 0.1645 G12EC ResNet 0.0129 0.0255 0.0286 0.0482 0.0566 0.0743 Transformer 0.0010 0.0205 0.0205 0.0558 0.0766 0.0947 LR -0.0083 0.0269 0.0282 0.0383 0.0223 0.0443 XGBoost 0.0395 0.0313 0.0504 0.0958 0.0828 0.1267 PTB and PTB-XL ResNet 0.0529 0.0263 0.0591 0.0805 0.0477 0.0936 Transformer 0.0336 0.0326 0.0468 0.0960 0.0603 0.1133 LR 0.0352 0.0448 0.0570 0.1041 0.0808 0.1318 XGBoost 0.0834 0.0291 0.0883 0.1496 0.0528 0.1586 SPH ResNet 0.0707 0.0160 0.0725 0.0682 0.0158 0.0700 Transformer 0.0772 0.0119 0.0781 0.1339 0.0230 0.1358 LR 0.0845 0.0364 0.0920 0.1256 0.0493 0.1349 XGBoost 0.0772 0.0156 0.0788 0.0946 0.0644 0.1144 LR = Logistic regression
4.2 Multi-source experiments
The results from the multi-source experiments, where data from multiple sources were used for training and validation are summarized in Figure 4. The results are reported per test set. The results of the 4-fold CV were higher than the LSO CV for both macro- and micro-AUC regardless of the model, as expected. The 4-fold CV AUCs are very similar regardless of the test source for each model, and variance over the folds is small. In contrast, there is more variability between LSO CV results for different test sources and also between the folds. The LSO CV estimates for both macro- and micro-AUC tend to be closer to the test AUCs than the 4-fold CV estimates. The ResNet and Transformer networks consistently achieve higher AUCs for each test source compared to LR and XGBoost, particularly in the LSO CV.
@@ 4.2 Multi-source experiments
The comparison between the errors of 4-fold CV and LSO CV estimates is presented in Table 5 for each model. The mean errors for the LSO CV are close to zero, indicating unbiased performance estimates In contrast, the 4-fold CV results mean errors demonstrate clear optimistic bias. On the other hand, the LSO CV has higher SD for the errors. Combining these two components of error together, LSO CV has lower RMSE, providing overall more reliable performance estimates in the experiments than 4-fold CV.
@@
ResNet Transformer LR XGBoost 4-fold LSO 4-fold LSO 4-fold LSO 4-fold LSO Macro- AUC Mean 0.1121 -0.0061 0.0403 -0.0299 0.0468 0.0005 0.0684 0.0043 SD 0.1128 0.1372 0.0256 0.0336 0.0409 0.0446 0.0330 0.0260 RMSE 0.1590 0.1373 0.0477 0.0450 0.0622 0.0446 0.0759 0.0264 Micro- AUC Mean 0.1261 0.0012 0.0621 -0.0215 0.1800 0.0084 0.0884 -0.0049 SD 0.0605 0.0812 0.0427 0.0583 0.0937 0.1441 0.0963 0.1033 RMSE 0.1399 0.0812 0.0754 0.0621 0.2029 0.1443 0.1307 0.1034 LR = Logistic regression
4.3 Source prediction experiment
The confusion matrices per input set are visualized in Figure 5. The corresponding accuracies for the input sets were 1) 0.969, 2) 0.973, and 3) 0.625. A trivial baseline model that always predicts the largest class (Chapman-Shaoxing and Ningbo) achieves 0.424 accuracy. Clearly, the model that uses only diagnoses already performs higher than this leveraging systematic differences in label distributions. However, much higher performance still can be achieved when making predictions based on ECG, age and sex. Incorporating diagnoses as additional features to this model did not lead to noticeable improvement in accuracy.
Figure 5: Confusion matrices for each input set in the multiclass classification task where the output is the data source of ECGs
The confusion matrices indicate that the Chapman-Shaoxing and Ningbo source was more frequently misclassified, and other sources were often misclassified as this source. This occurs especially when using only diagnoses as an input: The ECGs from the G12EC source are more often misclassified as originating from the Chapman-Shaoxing and Ningbo source than they are correctly classified. Similarly, ECGs from the PTB and PTB-XL source is often misclassified as the SPH data source.